Find the surface area of the part of the sphere

x^2+y^2+z^2=9

lying above the ellipse

x^2/4 + y^2/9 =1


1
Expert's answer
2021-05-18T18:11:50-0400
"S=\int\int_D\sqrt{1+(\partial z/\partial x)^2+(\partial z/\partial y)^2}dxdy"

"x^2+y^2+z^2=9"

"z^2=9-x^2-y^2"

"2z(\partial z/\partial x)=-2x"

"\partial z/\partial x=-x/z"

"2z(\partial z/\partial y)=-2y"

"\partial z/\partial y=-y/z"

"1+(\partial z/\partial x)^2+(\partial z/\partial y)^2=1+\dfrac{x^2}{z^2}+\dfrac{y^2}{z^2}"

"=\dfrac{z^2+x^2+y^2}{z^2}=\dfrac{9}{z^2}"

"z^2=9-x^2-y^2, z\geq0"

"S=\int\int_D\sqrt{1+(\partial z/\partial x)^2+(\partial z/\partial y)^2}dxdy"

"=\int\int_D\dfrac{3}{\sqrt{9-x^2-y^2}}dxdy"

"\dfrac{x^2}{4}+\dfrac{y^2}{9}\leq1"

"x=2r\cos\theta, y=3r\sin\theta"

"\dfrac{\partial(x,y)}{\partial(r,\theta)}=\begin{vmatrix}
 2\cos\theta & -2r\sin\theta \\
 3\sin\theta & 3r\cos\theta
\end{vmatrix}=6r"

"S=\int\int_G\dfrac{6r}{\sqrt{9-4r^2\cos^2\theta-9r^2\sin^2\theta}}drd\theta"


"=\displaystyle\int_{0}^{2\pi}\displaystyle\int_{0}^{1}\dfrac{6r}{\sqrt{9-4r^2-5r^2\sin^2\theta}}drd\theta"

"=6\displaystyle\int_{0}^{2\pi}\bigg(\dfrac{3-\sqrt{5}\sin\theta}{4+5\cos^2\theta}\bigg)d\theta"

"=6\pi"

"S=6\pi" square units